OpenStudy (anonymous):

A cone with a radius of 3" has a total area of 24 sq. in. Find the volume of the cone. 15pi 12pi

OpenStudy (agl202):

Need help?

OpenStudy (anonymous):

yep

OpenStudy (agl202):

So this is how u do it: l=5 r=3 h=? In the right angled triangle inside the cone that @DarthTony had drawn..we see that hypotenuse is l or s(in the figure he calls slant height or l as s) Base=r Height=h Using Pythagoras' theorem we get l2=h2+r2 or h2=l2−r2 solving this(putting values of l and r and then taking square root) we shall get h to be equal to 4. h=4 Hence volume= 13πr2h Putting values we get volume as=12pi in^3 (Answer C) Did you get it?

OpenStudy (anonymous):

yep thanks a lot:))

OpenStudy (agl202):

yep! No problem.